HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sep. 1, 2023

ELECTRON SPIN-TORSION AND SPIN-ROTATION EFFECTS IN THE OPEN-SHELL CH_{3}CO RADICAL^{1}

L. H. COUDERT, O. PIRALI, R. CHAHBAZIAN, M.-A. MARTIN-DRUMEL, Institut des Sciences Moléculaires d’Orsay, Université Paris-Saclay, Orsay, France; L. ZOU, R. A. MOTIYENKO, L. MARGULÈS, UMR 8523 - PhLAM - Physique des Lasers Atomes et Molécules, University of Lille, CNRS, F-59000 Lille, France

Although open-shell molecules displaying a fine interaction and non-rigid molecules displaying a large amplitude motion have been the subject of numerous spectroscopic investigations, fewer results are available about molecules displaying both a fine interaction and a large amplitude motion. A non-exhaustive list of such species include the open shell CH_{2}, CH_{2}OH, and Na_{3} species displaying^{2} respectively a large amplitude bending, torsion, and pseudorotation. The acetyl radical CH_{3}CO is a benchmark molecule for this subclass of molecules as it presents a low barrier internal rotation of its methyl group and a fine interaction. In the first spectroscopic investigation of the radical,^{3} a few low J, K_{a} = 0 transitions were measured in its microwave spectrum and a new theoretical model was developed to model its spectrum.

In this poster, new measurements carried out in the submillimeter wave spectrum of CH_{3}CO will be presented. More than 300 transitions have been assigned up to N = 22 and K_{a} = 5 in the ground torsional state for both A and E torsional species and for both F_{1} and F_{2} electron spin-rotation components. Based on a newly developed theoretical approach, a line frequency analysis has been carried out and line frequencies were reproduced with a unitless standard deviation of 1.2.

The analysis shows that two effects, ignored in the pioneering approach of Hirota et al.,^{3} are quite important. These are the electron spin-torsion coupling, described by a term in p_{α}S_{z}, where p_{α} is the momentum conjugated to the torsional angle α, and a dependence of the electron spin-rotation coupling tensor on this torsional angle. In the poster the new measurements will be presented and the spectroscopic parameters will be reported. These include the usual torsion-rotation parameters A, B, C, D_{ab}, ρ, and V_{3}, and the new fine interaction-torsion parameters.

1Supported by the Programme National “Physique et Chimie du Milieu Interstellaire” (PCMI) of CNRS/INSU with INC/INP co-funded by CEA and CNES

2Ohashi, Tsuura, Hougen, Ernst, and Rakowsky, J. Mol. Spec. 184 (1997) 22; Coudert, J. Chem.

Phys. 153 (2020) 144115; and Coudert, Chitarra, Spaniol, Loison, Martin-Drumel, and Pirali, J. Chem.

Phys. 156 (2022) 244301

3Hirota, Mizoguchi, Ohsima, Katoh, Sumiyoshi, and Endo, Mol. Phys. 105 (2007) 455 p-number: p066  Submitted on Sun, 11 Jun 2023 07:03:21 +0200